An advanced computer sends information to its various parts via infrared light pulses traveling through silicon fibers (n=3.50). To acquire data from memory, the central processing unit sends a light-pulse request to the memory unit. The memory unit processes the request, then sends a data pulse back to the central processing unit. The memory unit takes 0.52 ns to process a request.

If the information has to be obtained from memory in 2.10 ns , what is the maximum distance the memory unit can be from the central processing unit?
Answer #1

Solution :-

speed of light = 3x 10^8 m/s /3.5= 8.57 x 10^7 m/s

total time light travels= 2.10-0.52 = 1.58 ns

one way distance it travels 1.58 ns/2=0.79 ns= 0.79 x 10^-9 s

maximum distance = 8.57 x 10^7 x 0.79 x 10^-9 = 6.77 x 10^-2 m =6.77cm